OpenStudy (anonymous):

Find the indicated critical z value. Find zα/2 for α = 0.08. 1.96 2.65 1.41 1.75

jimthompson5910 (jim_thompson5910):

do you have a TI83 or TI84 calculator?

OpenStudy (anonymous):

i do not

jimthompson5910 (jim_thompson5910):

ok let me find another calculator for you

OpenStudy (anonymous):

okay thank you

jimthompson5910 (jim_thompson5910):

ok go here http://onlinestatbook.com/2/calculators/inverse_normal_dist.html type in "0.04" in the "area" box without quotes leave "Mean" and "SD" alone tell me what you get

jimthompson5910 (jim_thompson5910):

btw, I got 0.04 from dividing alpha = 0.08 in half

OpenStudy (anonymous):

okay I got 1.751

jimthompson5910 (jim_thompson5910):

me too

OpenStudy (anonymous):

can you help me with this? Assume that women's heights are normally distributed with a mean of 63.6 inches and a standard deviation of 2.5 inches. If 90 women are randomly selected, find the probability that they have a mean height between 62.9 inches and 64.0 inches. 0.1739 0.0424 0.7248 0.9318

OpenStudy (anonymous):

I'm horrible with standard deviations

jimthompson5910 (jim_thompson5910):

First we need the z-score for the rawscore xbar = 62.9 \[\LARGE z = \frac{\bar{x}-\mu}{\frac{\sigma}{\sqrt{n}}}\] \[\LARGE z = \frac{62.9-63.6}{\frac{2.5}{\sqrt{90}}}\] \[\LARGE z = ??\]

OpenStudy (anonymous):

I got -2.6563132345

OpenStudy (anonymous):

I think I'm wrong :(

jimthompson5910 (jim_thompson5910):

that's correct

jimthompson5910 (jim_thompson5910):

what's the z-score for the rawscore of xbar = 64.0 ?

OpenStudy (anonymous):

1.5178932769

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so P(62.9 < xbar < 64.0) = P(-2.66 < Z < 1.52)

jimthompson5910 (jim_thompson5910):

now you'll use a table or a calculator to compute P(-2.66 < Z < 1.52)

jimthompson5910 (jim_thompson5910):

do you know how to do that? or no?

OpenStudy (anonymous):

no :(

jimthompson5910 (jim_thompson5910):

do you have a table with you?

jimthompson5910 (jim_thompson5910):

a stats table that provides area under the standard normal curve

OpenStudy (anonymous):

no all I have is the z score table

jimthompson5910 (jim_thompson5910):

ok I'll be right back in a few minutes and show you how to use a table to answer this question. Then we can also go over using a calculator as well

OpenStudy (anonymous):

okay thank you so much. my teacher doesn't help me nearly as much.

jimthompson5910 (jim_thompson5910):

ok I'm back

jimthompson5910 (jim_thompson5910):

the z score table you have, is that accessible online? If so, can I get the link? if not, then I'll find another table we can use

jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

first let's find P(Z < -2.66)

jimthompson5910 (jim_thompson5910):

locate the row that starts with -2.6 then find the column that has 0.06 the row and column combine to form -2.66 the value in that row/column combo is the approximate area under the curve to the left of z = -2.66

jimthompson5910 (jim_thompson5910):

see attached to see what I mean

OpenStudy (anonymous):

.00391 ?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

what is P(Z < 1.52) equal to?

OpenStudy (anonymous):

.93574

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

subtract the two values to get the area between z = -2.66 and z = 1.52

jimthompson5910 (jim_thompson5910):

[ P(Z < 1.52) ] - [ P(Z < -2.66) ] [ 0.93574 ] - [ 0.00391 ] 0.93183

jimthompson5910 (jim_thompson5910):

That means P(-2.66 < Z < 1.52) = 0.93183 and P(62.9 < xbar < 64.0) = 0.93183

OpenStudy (anonymous):

it seems so much better now thanks

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

would this be a continuation of the first one? Find the critical value zα/2 that corresponds to a 98% confidence level. 2.575 1.75 2.33 2.05

jimthompson5910 (jim_thompson5910):

Confidence level = 98% alpha = 1-(confidence level) alpha = 1-0.98 alpha = 0.02

jimthompson5910 (jim_thompson5910):

alpha = 0.02 alpha/2 = 0.02/2 alpha/2 = 0.01

jimthompson5910 (jim_thompson5910):

you will use that calculator I gave you and type in 0.01 (where you typed in 0.04)

OpenStudy (anonymous):

i got 2.23 so 2.33

OpenStudy (anonymous):

2.32 I mean

OpenStudy (anonymous):

2.327

jimthompson5910 (jim_thompson5910):

which becomes 2.33, yep

OpenStudy (anonymous):

This is like the other one we just did right? For women aged 18–24, systolic blood pressures (in mm Hg) are normally distributed with a mean of 114.8 and a standard deviation of 13.1. If 23 women aged 18–24 are randomly selected, find the probability that their mean systolic blood pressure is between 119 and 122. 0.3343 0.9341 0.0833 0.0577

jimthompson5910 (jim_thompson5910):

yes just with different numbers

jimthompson5910 (jim_thompson5910):

rawscore: xbar = 119 z-score: z = ???

OpenStudy (anonymous):

I did the whole thing and got it! in the end I got 0.93407

OpenStudy (anonymous):

But I don't know how to do this one with no mean or standard deviation provided Can you help me? Use the given degree of confidence and sample data to construct a confidence interval for the population proportion p. A survey of 865 voters in one state reveals that 408 favor approval of an issue before the legislature. Construct the 95% confidence interval for the true proportion of all voters in the state who favor approval. 0.438 < p < 0.505 0.471 < p < 0.472 0.435 < p < 0.508 0.444 < p < 0.500

jimthompson5910 (jim_thompson5910):

Use the formulas: \[\Large L = \hat{p} - z*\sqrt{\frac{\hat{p}*(1-\hat{p})}{n}}\] \[\Large U = \hat{p} + z*\sqrt{\frac{\hat{p}*(1-\hat{p})}{n}}\] where... L = Lower endpoint of confidence interval U = Upper endpoint of confidence interval The confidence interval is of the form (L,U) which can be written as L < p < U \(\Large \hat{p}\) is the sample proportion (the symbol is read has "p hat") z = critical value based on the confidence level n = sample size

OpenStudy (anonymous):

okay thanks trying it now

OpenStudy (anonymous):

Okay I'm still a little confused can you help me with one?

jimthompson5910 (jim_thompson5910):

how far did you get

OpenStudy (anonymous):

well I'm trying to figure out what to substitute

jimthompson5910 (jim_thompson5910):

\[\Large \hat{p} = \frac{x}{n} = \frac{408}{865} \approx 0.471676 \approx 0.472\] x = number of successes n = sample size

jimthompson5910 (jim_thompson5910):

z = 1.96 approximately based off the 95% confidence level

jimthompson5910 (jim_thompson5910):

since P(-1.96 < Z < 1.96) = 0.95 approximately

OpenStudy (anonymous):

okay I get it. One second

OpenStudy (anonymous):

i keep getting -0.025 for the lower

jimthompson5910 (jim_thompson5910):

what are you getting for the value of \[\Large \sqrt{\frac{\hat{p}*(1-\hat{p})}{n}}\]

OpenStudy (anonymous):

0.017

jimthompson5910 (jim_thompson5910):

\[\Large L = \hat{p} - z*\sqrt{\frac{\hat{p}*(1-\hat{p})}{n}}\] \[\Large L = 0.472 - 1.96*\sqrt{\frac{0.472*(1-0.472)}{865}}\] \[\Large L = 0.472 - 1.96*0.017\] \[\Large L = 0.43868\] things will be a bit off due to rounding errors, but it's close enough

OpenStudy (anonymous):

Oh I see. it was my order of operations. I subtracted first.

OpenStudy (anonymous):

if i'm doing a problem and I get a z score higher than 3.9 does that mean I did it wrong?

jimthompson5910 (jim_thompson5910):

I would have to see the full problem

OpenStudy (anonymous):

Suppose the number of students in a class for the Business Statistics program at a University has a mean of 23 with a standard deviation of 4.3. If 15 classes are selected randomly, find the probability that the mean number of students is between 20 and 30. 0.9699 0.6824 0.8412 0.9966

jimthompson5910 (jim_thompson5910):

what z-scores did you get?

OpenStudy (anonymous):

-2.70 and 6.30

jimthompson5910 (jim_thompson5910):

me too

jimthompson5910 (jim_thompson5910):

so you need to compute P(-2.70 < Z < 6.30)

jimthompson5910 (jim_thompson5910):

P(-2.70 < Z < 6.30) = P(Z < 6.30) - P(Z < -2.70) = ?? - ??

jimthompson5910 (jim_thompson5910):

unfortunately, the table you have doesn't go up to 6.30

jimthompson5910 (jim_thompson5910):

so use something like wolfram alpha http://www.wolframalpha.com/input/?i=P(-2.70%20%3C%20Z%20%3C%206.30)&t=crmtb01&f=rc

OpenStudy (anonymous):

okay thanks.

OpenStudy (anonymous):

Assume that a sample is used to estimate a population proportion p. Find the margin of error E that corresponds to the given statistics and confidence level. Round the margin of error to four decimal places. 95% confidence; the sample size is 10,000, of which 40% are successes 0.0110 0.0072 0.0096 0.0126

OpenStudy (anonymous):

should I use the equation from earlier?

jimthompson5910 (jim_thompson5910):

\[\Large E = z*\sqrt{\frac{\hat{p}*(1-\hat{p})}{n}}\]

jimthompson5910 (jim_thompson5910):

z = 1.96 at the 95% CL http://www.wolframalpha.com/input/?i=95%25+Confidence \(\Large \hat{p} = 0.4\) n = 10,000

OpenStudy (anonymous):

i got 0.0096

jimthompson5910 (jim_thompson5910):

I did also

OpenStudy (anonymous):

Would I do this like the other? In one region, the September energy consumption levels for single-family homes are found to be normally distributed with a mean of 1050 kWh and a standard deviation of 218 kWh. If 50 different homes are randomly selected, find the probability that their mean energy consumption level for September is greater than 1075 kWh. 0.4562 0.0438 0.2090 0.2910

jimthompson5910 (jim_thompson5910):

what's the z-score?

OpenStudy (anonymous):

i got 0.81

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so you need to compute P(Z > 0.81)

OpenStudy (anonymous):

0.2090?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

is this like any of the others?

OpenStudy (anonymous):

When 430 randomly selected light bulbs were tested in a laboratory, 224 lasted more than 500 hours. Find a point estimate of the proportion of all light bulbs that last more than 500 hours. 0.519 0.521 0.343 0.479

jimthompson5910 (jim_thompson5910):

\[\Large \hat{p} = \frac{x}{n} = \frac{224}{430} \approx 0.52093 \approx 0.521\] x = number of successes n = sample size